OpenStudy (anonymous):

the average rainfall for 5 days was 3inches per day. it rained 1 inch on the first day, 4.5 inches on the second day, 2.9 inches on the third day, and 4 inches on the fourth day. how many inches did it rain on the fifth day. please help! i will give medals!!

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Make an equation. \(\sf \dfrac{1 + 4.5 + 2.9 + 4 + x}{5}=3\)

OpenStudy (igreen):

Solve for 'x'. The mean over 5 days is 3. To find the mean, add up all the inches of rainfall and divide by how many days. This gives us that equation.

OpenStudy (igreen):

First, multiply 5 to both sides, can you do that?

OpenStudy (anonymous):

when you say multiply 5 by both sides do you mean by all the average numbers?

OpenStudy (anonymous):

like 1, 4.5, 2.9, and 4

OpenStudy (igreen):

\(\sf \dfrac{1 + 4.5 + 2.9 + 4 + x}{5} \cdot 5=3\cdot 5\) \(\sf \dfrac{1 + 4.5 + 2.9 + 4 + x}{\cancel 5} \cdot \cancel{5}=3\cdot 5\) \(\sf 1 + 4.5 + 2.9 + 4 + x = 3 \cdot 5\) Simplify the right side of the equation.

OpenStudy (anonymous):

my final answer before I posted the question was 2.6 inches so do you know what i did wrong?

OpenStudy (igreen):

2.6 is correct. \(\sf 1 + 4.5 + 2.9 + 4 + x = 15\) \(\sf 12.4 + x = 15\) \(\sf x = 2.6\)

OpenStudy (anonymous):

oh ok sweet. I thought you were saying that 3.5 was the answer. thanks for the help!

OpenStudy (igreen):

No problem.